A series RL circuit containing a 0.110 H inductor has a time constant of 0.150 s. What is the resistance in the circuit?
Answer #1

time constant=L/R

=>R=L/time constant

=0.110/0.150

=0.73 ohm
Answer #2

As time constant= L/R

0.15 = 0.11/R

R=0.7333